A commercial jet and a private airplane fly from Denver to Phoenix. It takes the commercial jet 1.6 hours for the flight, and it takes the private airplane 2.4 hours. The speed of the commercial jet is 180 miles per hour faster than the speed of the private airplane. Find the speed, in miles per hour, of both airplanes.
Answer by ikleyn(53765)

.
A commercial jet and a private airplane fly from Denver to Phoenix.
It takes the commercial jet 1.6 hours for the flight, and it takes the private airplane 2.4 hours.
The speed of the commercial jet is 180 miles per hour faster than the speed of the private airplane.
Find the speed, in miles per hour, of both airplanes.
~~~~~~~~~~~~~~~~~~~~~~~~~~~

Let x be the speed of the airplane, in miles per hour.

Then the speed of the jet is (x+180) mph.


The travel distance is the same, so we write this equation for distances

    2.4x = 1.6*(x+180)  miles


Simplify and find x

    2.4x = 1.6x + 288

    2.4x - 1.6x = 288

        0.8x    = 288

           x    = 288/0.8 = 360.


ANSWER. The airplane speed is 360 mph.  The jet speed is 360+180 = 540 mph.

Solved.